The percent of aluminum in the compound, Al2X3, is 18.56 %. What is the molar mass of element X ?



1
Expert's answer
2021-09-12T00:43:38-0400

Solution:

Al2X3

%Al = (2 × Molar mass of Al / Molar mass of Al2X3) × 100%

18.56% = (2 × 26.98 g Al mol-1 / Molar mass of Al2X3) × 100%

0.1856 = (53.96 / Molar mass of Al2X3)

Molar mass of Al2X3 = 53.96 / 0.1856 = 290.73


Molar mass of Al2X3 = 2 × Molar mass of Al + 3 × Molar mass of X

290.73 = 2 × 26.98 + 3 × Molar mass of X

290.73 = 53.96 + 3 × Molar mass of X

Molar mass of X = (290.73 - 53.96) / 3 = 78.92 g/mol


Answer: The molar mass of element X is 78.92 g/mol